Solve the system of equations and give the value of x:
4x + 5y =6
15y + 18 - 12x
Please show how you arrived at answer: Thank you

Answer by stanbon(75887)

4x + 5y =6
15y + 18 - 12x
-----------------------
Rearrange:
4x + 5y = 6
12x+15y = -18
---------------------
Modify to get:
12x + 15y = 18
12x + 15y = -18
-----
Those two equations are a contradiction.
Ans: No solution
=======================
Cheers,
Stan H.